# Imperial College London

### v1

@@ III Analysis of a Microwave Linear Analog Computer
We have introduced and modeled the concept of MiLAC, highlighting that it consists of a reconfigurable microwave network with multiple ports where input signals are applied and output signals are read. In this section, we characterize the output $\mathbf{v}$ as a function of the input $\mathbf{u}$ and the admittance matrix of the microwave network $\mathbf{Y}$.
We begin by noticing that a MiLAC is described by the system of linear equations (1), (2), and (3), given by
@@
which fully characterizes the MiLAC in Fig. 1. To solve (6), we rewrite it in a more compact form as
@@
    If $\mathbf{A}$ is invertible, it holds
@@
    If $\mathbf{D}$ is invertible, it holds
@@
The proposition follows from [20, Theorem 2.1]. ∎
@@
Remarkably, (21)-(22) and (23)-(24) contain the expressions of the output vectors $\mathbf{v}_{1}$ and $\mathbf{v}_{2}$ as a function of the input $\mathbf{u}$ and the blocks of $\mathbf{P}$, which are directly related to the blocks of $\mathbf{Y}$ by (20).
@@
Note that (21)-(22) and (23)-(24) would require computationally expensive matrix-matrix products and matrix inversion operations if computed through a digital computer. However, they can be quickly computed by a MiLAC as the EM signal propagates within the microwave network at light speed. In other words, after properly setting the input $\mathbf{u}$ and the tunable components $\{Y_{i,k}\}_{i,k=1}^{P}$, (21)-(22) and (23)-(24) can be computed with complexity $\mathcal{O}(1)$, independently from the input size $N$ and output size $P=N+M$. Remarkably, the expressions in (22) and (24) are particularly relevant for computing the LMMSE estimator, as it is clarified in Section IV.
@@
MiLAC output LMMSE estimator
$\mathbf{v}_{2}=\left(\left(\mathbf{P}_{21}\mathbf{P}_{11}^{-1}\mathbf{P}_{12}-\mathbf{P}_{22}\right)^{-1}\mathbf{P}_{21}\mathbf{P}_{11}^{-1}\right)\mathbf{u}$ $\hat{\mathbf{x}}_{\textnormal{LMMSE},1}=\left(\left(\mathbf{H}^{H}\mathbf{C}_{\mathbf{n}}^{-1}\mathbf{H}+\mathbf{C}_{\mathbf{x}}^{-1}\right)^{-1}\mathbf{H}^{H}\mathbf{C}_{\mathbf{n}}^{-1}\right)\mathbf{y}$
$\mathbf{v}_{2}=\left(\mathbf{P}_{22}^{-1}\mathbf{P}_{21}\left(\mathbf{P}_{12}\mathbf{P}_{22}^{-1}\mathbf{P}_{21}-\mathbf{P}_{11}\right)^{-1}\right)\mathbf{u}$ $\hat{\mathbf{x}}_{\textnormal{LMMSE},2}=\left(\mathbf{C}_{\mathbf{x}}\mathbf{H}^{H}\left(\mathbf{H}\mathbf{C}_{\mathbf{x}}\mathbf{H}^{H}+\mathbf{C}_{\mathbf{n}}\right)^{-1}\right)\mathbf{y}$
IV Analog Computing for the LMMSE Estimator
@@
We have introduced a MiLAC as an analog computer that linearly processes microwave signals and we characterized its output as a function of its input and microwave network admittance matrix. In this section, we show that a MiLAC can be used to compute the LMMSE estimator for linear observation processes with significantly reduced computational complexity.
@@ IV-A LMMSE Estimator for Linear Observation Processes
where $\mathbf{H}\in\mathbb{C}^{Y\times X}$ is a known constant matrix, $\mathbf{x}\in\mathbb{C}^{X\times 1}$ is the unknown random vector with mean $\bar{\mathbf{x}}=\mathbf{0}_{X\times 1}$ and covariance matrix $\mathbf{C}_{\mathbf{x}}$, and $\mathbf{n}\in\mathbb{C}^{Y\times 1}$ is the random noise vector with mean $\bar{\mathbf{n}}=\mathbf{0}_{Y\times 1}$ and covariance matrix $\mathbf{C}_{\mathbf{n}}$. According to estimation theory, the LMMSE estimator $\hat{\mathbf{x}}_{\text{LMMSE}}$ of $\mathbf{x}$ is defined as the linear function of $\mathbf{y}$ minimizing the mean square error (MSE) $\text{E}[\|\hat{\mathbf{x}}_{\text{LMMSE}}-\mathbf{x}\|_{2}^{2}]$ [21, Chapter 2]. The expression of the LMMSE estimator for a linear observation process is available in closed form, as given by the following proposition.
@@
$\hat{\mathbf{x}}_{\textnormal{LMMSE},2}$ $=\left(\mathbf{C}_{\mathbf{x}}\mathbf{H}^{H}\left(\mathbf{H}\mathbf{C}_{\mathbf{x}}\mathbf{H}^{H}+\mathbf{C}_{\mathbf{n}}\right)^{-1}\right)\mathbf{y}.$ (29)
@@
The proposition follows from [21, Theorem 12.1]. ∎
Interestingly, there is a remarkable similarity between the LMMSE expressions in (28)-(29) and the expressions characterizing the output $\mathbf{v}_{2}$ of a MiLAC in (22) and (24), as highlighted in Tab. I. Thus, we can efficiently compute the LMMSE estimator for linear observation processes with a MiLAC, as specified in the following.
@@ IV-B Computing the LMMSE Estimator with a MiLAC
are two values of $\mathbf{P}_{\text{LMMSE}}$ giving the desired output. Note that this MiLAC has $N=Y$ input ports and $M=X$ output ports with no input, giving a total of $P=X+Y$ ports. Interestingly, analog computers have been proposed to compute generalized regression, ridge regression, and linear regression through resistive memory arrays [15, 16, 17], all operations that can be seen as special cases of the LMMSE estimator, confirming the generality of our model.
V Performance Evaluation

### v2

@@ III Analysis of a Microwave Linear Analog Computer
We have introduced and modeled the concept of MiLAC, highlighting that it consists of a reconfigurable microwave network with multiple ports where input signals are applied and output signals are read. In this section, we characterize the output $\mathbf{v}$ as a function of the input $\mathbf{u}$ and the admittance matrix of the microwave network $\mathbf{Y}$. For better clarity, we distinguish between the cases where the input is applied only to some ports ($N<P$) and to all ports ($N=P$).
III-A Analysis of a MiLAC with Input on Some Ports
In the case the input is applied only to some ports of the reconfigurable microwave network, i.e., $N<P$ and $M>0$, the MiLAC is described by the system of linear equations (2), (3), and (4), given by
@@ III-A Analysis of a MiLAC with Input on Some Ports
which characterizes the MiLAC in Fig. 1. To solve (8), we rewrite it in a more compact form as
@@
    If $\mathbf{A}$ is invertible, $\mathbf{P}$ invertible implies that $\mathbf{C}\mathbf{A}^{-1}\mathbf{B}-\mathbf{D}$ is invertible and
@@
    $\mathbf{B}^{\prime}=\mathbf{A}^{-1}\mathbf{B}\left(\mathbf{C}\mathbf{A}^{-1}\mathbf{B}-\mathbf{D}\right)^{-1},$ (18)
@@
    $\mathbf{D}^{\prime}=-\left(\mathbf{C}\mathbf{A}^{-1}\mathbf{B}-\mathbf{D}\right)^{-1}.$ (20)
@@
    If $\mathbf{D}$ is invertible, $\mathbf{P}$ invertible implies that $\mathbf{B}\mathbf{D}^{-1}\mathbf{C}-\mathbf{A}$ is invertible and
@@
    $\mathbf{B}^{\prime}=\left(\mathbf{B}\mathbf{D}^{-1}\mathbf{C}-\mathbf{A}\right)^{-1}\mathbf{B}\mathbf{D}^{-1},$ (22)
@@
    $\mathbf{D}^{\prime}=\mathbf{D}^{-1}-\mathbf{D}^{-1}\mathbf{C}\left(\mathbf{B}\mathbf{D}^{-1}\mathbf{C}-\mathbf{A}\right)^{-1}\mathbf{B}\mathbf{D}^{-1}.$ (24)
@@
The proposition follows from [38, Theorem 2.1]. ∎
@@
Remarkably, (27)-(28) and (29)-(30) contain the expressions of the output vectors $\mathbf{v}_{1}$ and $\mathbf{v}_{2}$ as a function of the input $\mathbf{u}$ and the blocks of $\mathbf{P}$, which are directly related to the blocks of $\mathbf{Y}$ by (26), when $M>0$.
@@
Remark 1.
Note that (27)-(28) and (29)-(30) would require computationally expensive matrix-matrix products and matrix inversion operations if computed through a digital computer. However, they can be quickly computed by a MiLAC as the EM signal propagates within the microwave network at light speed. In other words, after properly setting the input $\mathbf{u}$ and the tunable components $\{Y_{i,k}\}_{i,k=1}^{P}$, (27)-(28) and (29)-(30) can be computed with complexity $\mathcal{O}(1)$, independently from the input size $N$ and output size $P=N+M$. Remarkably, the expressions in (27)-(28) and (29)-(30) are particularly relevant for computing the LMMSE estimator and the covariance matrix of its error, as it will be clarified in Section IV.
Figure 4: Representation of an $N$-port MiLAC with input on all ports.
@@
MiLAC output LMMSE estimator and its error covariance matrix $\begin{bmatrix}\mathbf{P}_{11}&\mathbf{P}_{12}\\ \mathbf{P}_{21}&\mathbf{P}_{22}\end{bmatrix}$
$\mathbf{v}_{2}=\left(\left(\mathbf{P}_{21}\mathbf{P}_{11}^{-1}\mathbf{P}_{12}-\mathbf{P}_{22}\right)^{-1}\mathbf{P}_{21}\mathbf{P}_{11}^{-1}\right)\mathbf{u}$ $\hat{\mathbf{x}}_{\textnormal{LMMSE},1}=\left(\left(\mathbf{H}^{H}\mathbf{C}_{\mathbf{n}}^{-1}\mathbf{H}+\mathbf{C}_{\mathbf{x}}^{-1}\right)^{-1}\mathbf{H}^{H}\mathbf{C}_{\mathbf{n}}^{-1}\right)\mathbf{y}$ [±$\begin{bmatrix}\pm\mathbf{C}_{\mathbf{n}}&\mathbf{H}\\ \mathbf{H}^{H}&\mp\mathbf{C}_{\mathbf{x}}^{-1}\end{bmatrix}$
$\mathbf{v}_{2}=\left(\mathbf{P}_{22}^{-1}\mathbf{P}_{21}\left(\mathbf{P}_{12}\mathbf{P}_{22}^{-1}\mathbf{P}_{21}-\mathbf{P}_{11}\right)^{-1}\right)\mathbf{u}$ $\hat{\mathbf{x}}_{\textnormal{LMMSE},2}=\left(\mathbf{C}_{\mathbf{x}}\mathbf{H}^{H}\left(\mathbf{H}\mathbf{C}_{\mathbf{x}}\mathbf{H}^{H}+\mathbf{C}_{\mathbf{n}}\right)^{-1}\right)\mathbf{y}$ [±$\begin{bmatrix}\pm\mathbf{C}_{\mathbf{n}}&\mathbf{H}\\ \mathbf{H}^{H}&\mp\mathbf{C}_{\mathbf{x}}^{-1}\end{bmatrix}$
$\mathbf{v}_{1}=\left(\mathbf{P}_{11}^{-1}-\mathbf{P}_{11}^{-1}\mathbf{P}_{12}\left(\mathbf{P}_{21}\mathbf{P}_{11}^{-1}\mathbf{P}_{12}-\mathbf{P}_{22}\right)^{-1}\mathbf{P}_{21}\mathbf{P}_{11}^{-1}\right)\mathbf{u}$ $\mathbf{C}_{\mathbf{e},1}=\mathbf{C}_{\mathbf{x}}-\mathbf{C}_{\mathbf{x}}\mathbf{H}^{H}\left(\mathbf{H}\mathbf{C}_{\mathbf{x}}\mathbf{H}^{H}+\mathbf{C}_{\mathbf{n}}\right)^{-1}\mathbf{H}\mathbf{C}_{\mathbf{x}}$ $\begin{bmatrix}\mathbf{C}_{\mathbf{x}}^{-1}&\mathbf{H}^{H}\\ \mathbf{H}&-\mathbf{C}_{\mathbf{n}}\end{bmatrix}$
$\mathbf{v}_{1}=-\left(\mathbf{P}_{12}\mathbf{P}_{22}^{-1}\mathbf{P}_{21}-\mathbf{P}_{11}\right)^{-1}\mathbf{u}$ $\mathbf{C}_{\mathbf{e},2}=\left(\mathbf{H}^{H}\mathbf{C}_{\mathbf{n}}^{-1}\mathbf{H}+\mathbf{C}_{\mathbf{x}}^{-1}\right)^{-1}$ $\begin{bmatrix}\mathbf{C}_{\mathbf{x}}^{-1}&\mathbf{H}^{H}\\ \mathbf{H}&-\mathbf{C}_{\mathbf{n}}\end{bmatrix}$
III-B Analysis of a MiLAC with Input on All Ports
In the case the input is applied on all the ports of the reconfigurable microwave network, i.e., $N=P$ and $M=0$, we have $\mathbf{v}=\mathbf{v}_{1}$ and $\mathbf{i}=\mathbf{i}_{1}$, as represented in Fig. 4. Thus, this MiLAC is characterized by the system
$\begin{cases}\mathbf{i}_{1}&=\mathbf{Y}\mathbf{v}_{1}\\ \mathbf{i}_{1}&=Y_{0}\left(\mathbf{u}-\mathbf{v}_{1}\right)\end{cases},$ (33)
which yields
$\mathbf{Y}\mathbf{v}_{1}=Y_{0}\left(\mathbf{u}-\mathbf{v}_{1}\right),$ (34)
and allows us to express the output $\mathbf{v}_{1}$ as
$\mathbf{v}_{1}=\mathbf{P}^{-1}\mathbf{u},$ (35)
where
$\mathbf{P}=\frac{\mathbf{Y}}{Y_{0}}+\mathbf{I}_{N},$ (36)
giving the output vector $\mathbf{v}_{1}$ as a function of the input $\mathbf{u}$ and the matrix $\mathbf{Y}$, when $M=0$. Assuming that the tunable components $\{Y_{i,k}\}_{i,k=1}^{N}$ can take any arbitrary complex value, a MiLAC can compute (35) given any matrix $\mathbf{P}$. To this end, we first compute $\mathbf{Y}$ from any arbitrary $\mathbf{P}$ as
$\mathbf{Y}=Y_{0}\mathbf{P}-Y_{0}\mathbf{I}_{N},$ (37)
which follows from (36), and then we set the tunable components $\{Y_{i,k}\}_{i,k=1}^{N}$ as a function of $\mathbf{Y}$ as in (7). By substituting (37) into (7), we obtain the expression of the tunable admittance components as a function of any arbitrary $\mathbf{P}$ as
$Y_{i,k}=\begin{cases}-Y_{0}\left[\mathbf{P}\right]_{i,k}&i\neq k\\ Y_{0}\sum_{n=1}^{N}\left[\mathbf{P}\right]_{n,k}-Y_{0}&i=k\end{cases},$ (38)
for $i,k=1,\ldots,N$, enabling the MiLAC to compute (35) for any given $\mathbf{P}$.
Remark 2.
Computing (35) with a digital computer would involve the expensive computation of the matrix inverse $\mathbf{P}^{-1}$. However, (35) can be computed instantly by a MiLAC, regardless of the number of inputs $N$ and without requiring any matrix inversion operation. Remarkably, this enables an efficient computation of the matrix inverse, as it will be detailed in Section IV.
III-C Primitive Operations of a MiLAC
In light of the analysis carried out in Sections III-A and III-B, a MiLAC can be used as a computing device able to perform computing operations in the analog domain. These operations are computed as the EM signal propagates within the microwave network at light speed and can be read as output on the network ports. As a digital computer can compute a set of primitive operations contained in its instruction set architecture (ISA), also a MiLAC can compute a set of primitive operations which can be sequentially used to evaluate more complicated functions. We identify three primitive operations that a MiLAC with $P$ ports can compute as a function of its input $\mathbf{u}$ applied on $N$ ports and the tunable admittance components $\{Y_{i,k}\}_{i,k=1}^{P}$. The first primitive, denoted as
$\mathbf{v}_{1}=\text{MiLAC}_{1}\left(\mathbf{u},\{Y_{i,k}\}_{i,k=1}^{P}\right),$ (39)
returns in $\mathbf{v}_{1}$ the output on the first $N$ ports of the MiLAC. Mathematically, this primitive is expressed as (27) or (29) depending on the invertibility of $\mathbf{P}_{11}$ and $\mathbf{P}_{22}$, where $\mathbf{P}$ is defined as in (12). The second primitive operation, denoted as
$\mathbf{v}_{2}=\text{MiLAC}_{2}\left(\mathbf{u},\{Y_{i,k}\}_{i,k=1}^{P}\right),$ (40)
gives in $\mathbf{v}_{2}$ the output on the last $M$ ports of the MiLAC, with $M=P-N$. This primitive is mathematically given by (28) or (30) depending on the invertibility of $\mathbf{P}_{11}$ and $\mathbf{P}_{22}$. Last, the third primitive is denoted as
$\mathbf{v}=\text{MiLAC}_{3}\left(\mathbf{u},\{Y_{i,k}\}_{i,k=1}^{P}\right),$ (41)
and returns the output on all the $P$ ports of the MiLAC, i.e., the concatenation of (39) and (40). The primitive (41) coincides with (39) when the input is applied on all the ports, i.e., $N=P$ and $M=0$. The sequential use of different primitive operations computed in the analog domain can enable the efficient analog computation of complex algorithms. Nevertheless, it remains less investigated in literature [4]-[36]. In the following, we show how to exploit these three primitives together with low-complexity digital operations to efficiently compute functions with practical interest, namely the LMMSE estimator, the matrix inversion, and the Kalman filter.
IV Analog Computing for the LMMSE Estimator and Matrix Inversion
@@
We have introduced a MiLAC as an analog computer that linearly processes microwave signals and we characterized its output as a function of its input and microwave network admittance matrix. In this section, we show that a MiLAC can be used to compute the LMMSE estimator for linear observation processes and invert a matrix with significantly reduced computational complexity.
@@ IV-A LMMSE Estimator for Linear Observation Processes
where $\mathbf{H}\in\mathbb{C}^{Y\times X}$ is a known constant matrix, $\mathbf{x}\in\mathbb{C}^{X\times 1}$ is the unknown random vector with mean $\bar{\mathbf{x}}=\mathbf{0}_{X\times 1}$ and covariance matrix $\mathbf{C}_{\mathbf{x}}$, and $\mathbf{n}\in\mathbb{C}^{Y\times 1}$ is the random noise vector with mean $\bar{\mathbf{n}}=\mathbf{0}_{Y\times 1}$ and covariance matrix $\mathbf{C}_{\mathbf{n}}$. We assume that $\mathbf{C}_{\mathbf{x}}$ and $\mathbf{C}_{\mathbf{n}}$ are known and invertible, and that the cross-covariance matrix of $\mathbf{x}$ and $\mathbf{n}$ is $\mathbf{C}_{\mathbf{x}\mathbf{n}}=\mathbf{0}_{X\times Y}$. According to estimation theory, an estimator $\hat{\mathbf{x}}\in\mathbb{C}^{X\times 1}$ of $\mathbf{x}$ is any function of $\mathbf{y}$, its estimation error vector writes as $\mathbf{e}=\hat{\mathbf{x}}-\mathbf{x}$, and the mean square error (MSE) is given by $\text{E}[\|\hat{\mathbf{x}}-\mathbf{x}\|_{2}^{2}]$ [39, Chapter 2]. The LMMSE estimator $\hat{\mathbf{x}}_{\text{LMMSE}}$ of $\mathbf{x}$ is defined as the linear estimator minimizing the MSE, i.e.,
$\hat{\mathbf{x}}_{\text{LMMSE}}=\underset{\hat{\mathbf{x}}}{\text{min}}\;\text{E}\left[\left\|\hat{\mathbf{x}}-\mathbf{x}\right\|_{2}^{2}\right]\;\mathsf{\mathrm{s.t.}}\;\hat{\mathbf{x}}=\mathbf{W}\mathbf{y}+\mathbf{b},$ (43)
where $\mathbf{W}\in\mathbb{C}^{X\times Y}$ and $\mathbf{b}\in\mathbb{C}^{X\times 1}$. The expression of the LMMSE estimator for a linear observation process is available in closed form, as given by the following proposition.
@@
$\hat{\mathbf{x}}_{\textnormal{LMMSE},2}$ $=\left(\mathbf{C}_{\mathbf{x}}\mathbf{H}^{H}\left(\mathbf{H}\mathbf{C}_{\mathbf{x}}\mathbf{H}^{H}+\mathbf{C}_{\mathbf{n}}\right)^{-1}\right)\mathbf{y},$ (45)
where it holds that $\hat{\mathbf{x}}_{\textnormal{LMMSE},1}=\hat{\mathbf{x}}_{\textnormal{LMMSE},2}$. Furthermore, the LMMSE estimation error vector $\mathbf{e}=\hat{\mathbf{x}}_{\text{LMMSE}}-\mathbf{x}$ has mean $\bar{\mathbf{e}}=\mathbf{0}_{X\times 1}$ and covariance matrix $\mathbf{C}_{\mathbf{e}}$ equivalently given by
$\mathbf{C}_{\mathbf{e},1}$ $=\mathbf{C}_{\mathbf{x}}-\mathbf{C}_{\mathbf{x}}\mathbf{H}^{H}\left(\mathbf{H}\mathbf{C}_{\mathbf{x}}\mathbf{H}^{H}+\mathbf{C}_{\mathbf{n}}\right)^{-1}\mathbf{H}\mathbf{C}_{\mathbf{x}},$ (46)
$\mathbf{C}_{\mathbf{e},2}$ $=\left(\mathbf{H}^{H}\mathbf{C}_{\mathbf{n}}^{-1}\mathbf{H}+\mathbf{C}_{\mathbf{x}}^{-1}\right)^{-1},$ (47)
where we have $\mathbf{C}_{\mathbf{e},1}=\mathbf{C}_{\mathbf{e},2}$.
@@
The proposition follows from [39, Theorem 12.1]. ∎
Interestingly, there is a remarkable similarity between the LMMSE expressions in (44)-(47) and the expressions characterizing the output of a MiLAC, as highlighted in Tab. I. Specifically, we notice a similarity between the LMMSE estimator expressions in (44) and (45) and the expressions of $\mathbf{v}_{2}$ in (28) and (30), respectively. Likewise, (46) and (47) have the same structure as the matrix multiplying $\mathbf{u}$ in (27) and (29), respectively. Thus, we can efficiently compute the LMMSE estimator and the covariance matrix of its error with a MiLAC, as specified in the following.
@@ IV-B Computing the LMMSE Estimator with a MiLAC
are two values of $\mathbf{P}_{\text{LMMSE}}$ giving the desired output. Note that this MiLAC has $N=Y$ input ports and $M=X$ output ports with no input, giving a total of $P=X+Y$ ports. The steps necessary to compute the LMMSE estimator in the analog domain with a MiLAC are summarized in Alg. 1. Note that analog computers have been proposed to compute generalized regression, ridge regression, and linear regression through resistive memory arrays [29, 30, 35]. Interestingly, all these operations can be seen as special cases of the LMMSE estimator, confirming the generality of our model.
1:$\mathbf{y}$, $\mathbf{H}$, $\mathbf{C}_{\mathbf{x}}^{-1}$, and $\mathbf{C}_{\mathbf{n}}$.
2:$\hat{\mathbf{x}}_{\text{LMMSE}}$.
3: Set $\{Y_{i,k}\}_{i,k=1}^{P}$ by (32) with $\mathbf{P}=\mathbf{P}_{\text{LMMSE}}$.
4:$\hat{\mathbf{x}}_{\text{LMMSE}}=\text{MiLAC}_{2}(\mathbf{y},\{Y_{i,k}\}_{i,k=1}^{P})$.
Algorithm 1 LMMSE estimator with MiLAC
1:$\mathbf{H}$, $\mathbf{C}_{\mathbf{x}}^{-1}$, and $\mathbf{C}_{\mathbf{n}}$.
2:$\mathbf{C}_{\mathbf{e}}$.
3:Set $\{Y_{i,k}\}_{i,k=1}^{P}$ by (32) with $\mathbf{P}=\mathbf{P}_{\text{Cov}}$.
4:for $n=1\textbf{ to }N$ do
5:   $[\mathbf{C}_{\mathbf{e}}]_{:,n}=\text{MiLAC}_{1}([\mathbf{I}_{N}]_{:,n},\{Y_{i,k}\}_{i,k=1}^{P})$.
6:end for
Algorithm 2 Covariance matrix of the LMMSE estimator error with MiLAC
1:$\mathbf{P}$.
2:$\mathbf{P}^{-1}$.
3:Set $\{Y_{i,k}\}_{i,k=1}^{N}$ by (38).
4:for $n=1\textbf{ to }N$ do
5:   $[\mathbf{P}^{-1}]_{:,n}=\text{MiLAC}_{3}([\mathbf{I}_{N}]_{:,n},\{Y_{i,k}\}_{i,k=1}^{N})$.
6:end for
Algorithm 3 Matrix inversion with MiLAC
Figure 6: Computational complexity of inverting a matrix with size $N$.
Remark 3.